OpenStudy (anonymous):

The longer diagonal of a parallelogram measures 62 cm and makes an angle of 30 degrees with the base. Find the area of the parallelogram if the diagonals intersect at angle of 70 degrees.

OpenStudy (perl):



OpenStudy (perl):



OpenStudy (anonymous):

The four small triangles determined by the diagonals of a parallelogram have equal areas. So it is enough to find the area of one and multiply it by 4. Let us find the area of a triangle having one side a and 30 snd 70 angles that have side a. Let x be the the other side adjacent to 30 then \[ \frac {x}{\sin 70}=\frac {a}{\sin 80}\\ x= {\sin 70}\frac {a}{\sin 80}=0.954189 a \] Area of the triangle \[ \frac 12 x a \sin(30)=\frac 14 x a=\frac 1 4 0.954189 a^2=\frac 1 4 0.954189 \times31^2=229.244 \] Area of the paralleogram \[ 4 \times 229.244=916.976 \]